OpenStudy (anonymous):

About how accurately must be the exterior diameter of a 10 meters high cylindrical storage tank measured to calculate the amount of paint it will take to paint the side of the tank to within %5 of it's true value? Please explain thoroughly I have a test soon!

OpenStudy (anonymous):

im not sure about this one ...

OpenStudy (whpalmer4):

not sure why you closed the other question where I was trying to help you... we want the surface area just of the side of the cylinder, not the the ends, because the question says "paint the side of the tank", right? what is the formula for the area of the side of the cylinder?

OpenStudy (whpalmer4):

the area of the side of the cylinder depends on two factors, the height, and the radius (or diameter), right? it's just a rectangle rolled up so two edges meet...

OpenStudy (whpalmer4):

one side of the rectangle has the same length as the height of the cylinder. the other side has the same length as the circumference of the cylinder, right?

OpenStudy (anonymous):

is it pi(R^2)(h)

OpenStudy (whpalmer4):

@cvlibol no, that is the volume. think about the units you get with that formula: \[\pi * ([length])^2 * [length] = \pi* [length]^3\]